OpenStudy (anonymous):

If possible, find a solution to tan(4theta+1)=-6.

OpenStudy (anonymous):

\[\tan(4\theta+1) = -6\] \[4\theta+1 = \arctan(-6) \] \[\theta = {1\over4}(\arctan[-6]-1) \]